OpenStudy (anonymous):

In solving the equation (x + 1)(x – 2) = 4, Eric stated that the solution would be x + 1 = 4 => x = 3 or (x – 2) = 4 => x = 6 However, at least one of these solutions fails to work when substituted back into the original equation. Why is that? Please help Eric to understand better; solve the problem yourself, and explain your reasoning.

OpenStudy (anonymous):

to put it simplly (x+1)(x-2) must be equal to "0" not 4 for that method to work so that is y he is wrong

OpenStudy (anonymous):

what you must do is this (x+1)(x-2)=4 x^2-x-2=4 x^2-x-6=0 (x+2)(x-3)=0 therefore x=-2 or x=3

OpenStudy (anonymous):

understood?

OpenStudy (anonymous):

yes thanks

OpenStudy (anonymous):

:)